Solve:
y-1 = sqrt 2y-2

Answer by stanbon(75887)

y-1 = sqrt (2y-2)
----
(y-1)^2 = 2y-2
------
y^2 -2y + 1 = 2y - 2
----
y^2 - 4y + 3 = 0
-----
Factor::
(y-1)(y-3) = 0
----
y = 1 or y = 3
--------------
Cheers,
Stan H.
---------------